# To CoT or To Loop? A Formal Comparison Between Chain-of-Thought and Looped Transformers

### v1

@@ 1 Introduction
  • Does separation exist beyond the logarithmic regime? → Affirmative; see Section 4.
  • Are Looped Transformers universally more expressive than CoT? → Negative; see Section 5.
@@ 2 Preliminaries and Notation
We consider models based on Transformer blocks that are applied recursively. Specifically, we analyze two architectures that differ in how intermediate computations are structured and reused across steps. We assume saturated attention, where attention weights are uniformly assigned to the tokens with the highest scores. For the feedforward computation, we consider either a standard fully connected layer or a mixture-of-experts (MoE) layer (Shazeer et al., 2017; Csordás et al., 2024). Formal definitions of the Transformer block are deferred to Appendix A.
@@ 2.3 Notation
For any $n\in\mathbb{N}^{+}$, we write $[n]\coloneq\{1,2,\ldots,n\}$. For functions $f,g:\mathbb{N}\to\mathbb{N}$, we write $f(n)\in O(g(n))$ iff there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$. We denote by $\mathsf{poly}(n)$ the set of functions that grow at most polynomially, i.e., $\mathsf{poly}(n)\coloneq\left\{f:\mathbb{N}\to\mathbb{N}\;\middle|\;\exists k\in\mathbb{N},\;\exists c>0,\;\forall n\in\mathbb{N},\;f(n)\leq c\cdot n^{k}\right\}.$ Given two complexity classes ${\mathcal{C}}_{1}$ and ${\mathcal{C}}_{2}$, we write ${\mathcal{C}}_{1}\setminus{\mathcal{C}}_{2}$ to denote the set of decision problems (referred to as languages in computational complexity theory) that belong to ${\mathcal{C}}_{1}$ but not to ${\mathcal{C}}_{2}$.
@@ 3 General Problem Setting
This section begins by reviewing the problem settings considered in prior work, as summarized in Table 1, and identifying the remaining open questions. We then formalize the two problem settings addressed in this study, specifying their underlying assumptions and objectives.
@@ 3.3 Our Problem Settings
We consider two fundamental classes of computational problems: the deterministic problem and the probabilistic problem, elaborated in Section 4 and Section 5, respectively. This section provides high-level definitions that specify the input–output structure in each setting; precise formulations are deferred to the corresponding sections. These problem classes differ in several fundamental aspects, such as output uniqueness, the presence of a target distribution, and the inclusion of accuracy parameters.
@@ 4 Looped Transformer Enables Efficient Parallel Solutions
This section analyzes the fundamental computational differences between CoT and Looped TF in the deterministic setting. We formalize deterministic computations using computation graphs, as introduced in Section 4.1. In Section 4.2, we show how CoT and Looped TF differ in processing the same computation graphs, highlighting their respective strengths. Finally, Section 4.3 presents a separation in their computational complexity in the polylogarithmic regime.
@@ 4.1 Problem Setting
Let $\Sigma$ be a finite alphabet. A computation graph is a directed acyclic graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ that defines a function $G_{n}:\Sigma^{n}\to\Sigma^{m}$, where $\mathcal{F}_{n}=\{f_{n}:\Sigma^{\ell_{n}}\to\Sigma\mid\ell_{n}\in\mathbb{N}\}$ is a finite set of local functions associated with each node. The graph $G_{n}$ consists of (i) input nodes with in-degree 00, (ii) function nodes labeled by $f\in\mathcal{F}_{n}$ whose in-degree matches the arity of the corresponding function, and (iii) $m$ output nodes $O\subseteq V_{n}$ with out-degree 00. The overall function is obtained by evaluating the graph in topological order. The size of the graph is $|V_{n}|$, and its depth is the length of the longest path from an input node to an output node. We assume that the size of the computation graph $|V_{n}|$ is bounded by a polynomial in the input length $n$.
@@
To characterize the expressiveness of CoT and Looped TF, we analyze how each simulates computation graphs. We begin by formalizing the assumptions under which computations can be implemented by Transformer-based models (see Section B.1 for details).
@@
Remark. Proofs are given in Appendix B, with illustrations in Fig. 1. Informally, Theorem 4.4 shows that CoT simulates computations by decoding nodes in topological order, whereas Theorem 4.7 shows that Looped TF simulates computations in a layer-wise and parallel fashion. This contrast reflects complementary strengths: CoT employs intermediate steps as scratchpad memory, supporting general computation without structural constraints. In contrast, Looped TF exploits structural parallelism to achieve greater efficiency, but requires additional assumptions.
@@ 4.3 Separation between CoT and Looped Transformers in Polylogarithmic Time
Let $\mathsf{CoT}[T(n),d(n),p(n)]$ (resp. $\mathsf{LOOP}[T(n),d(n),p(n)]$) denote the set of languages :${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a CoT (resp. Looped TF) $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(p(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token, after $O(T(n))$ decode steps (resp. loops), is ${\mathcal{L}}(x)$, i.e., $M_{n}(x)={\mathcal{L}}(x),\text{ for all }x\in\{0,1\}^{n}.$
@@
Since Boolean circuits allow polynomial-size, we assume a polynomial embedding size. Prior work has shown that CoT is equivalent to Boolean circuits composed of $\mathsf{TC}^{0}$ gates, with circuit “size” proportional to the number of CoT steps. This class $\mathsf{SIZE}^{\mathsf{TC}^{0}}$ is defined in Definition B.6.
@@
There are several concrete problems in (nonuniform) $\mathsf{NC}^{2}$: solving linear equations, testing for perfect matching in bipartite graphs, and membership testing for fixed context-free grammars, denoted by LIN, MATCH, and FCFL, respectively (see Section B.6 for details). Then we have:
@@ 5 Chain-of-Thought Enables Probabilistic Reasoning
In this section, we investigate the expressiveness of CoT with stochastic decoding for probabilistic tasks. Section 5.1 introduces the task formulation and defines the model. Section 5.2 analyzes the expressive power of CoT in an idealized setting and establishes a formal separation from Looped TF. Finally, Section 5.3 shows that even weak CoT, when combined with self-consistency, exhibits sufficient expressive power.
@@ 5.1 Problem Setting
A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if any solution $y=y_{1}\dots y_{n}$ to input $x$ can be constructed incrementally, such that a prefix $y_{1:\sigma(x)}$ determines a subproblem $\psi(x,y_{1:\sigma(x)})$ whose solution completes $y$. See Definition C.4 for a formal definition.
@@
Approximate sampling is closely related to approximate counting (Jerrum et al., 1986). For a relation $R$, define $f(x)\coloneq\lvert R(x)\rvert$, the number of solutions $y$ such that $(x,y)\in R$. A Fully Polynomial-Time Approximation Scheme (FPTAS) is a deterministic algorithm that, for any input $x$ and $\varepsilon>0$, outputs $\hat{f}(x)$ such that $(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ in time polynomial in $|x|$ and $1/\varepsilon$. A Fully Polynomial Randomized Approximation Scheme (FPRAS) is a randomized version that additionally takes $\delta>0$, and satisfies the same guarantee with probability at least $1-\delta$. See Appendix C for details.
@@ 5.3 Expressiveness of Weak CoT
While Lemma 5.5 analyzes an idealized setting where a CoT, viewed as a PTM, is assumed to receive explicit error parameters, such assumptions diverge from the behavior of practical LLMs. In practice, models are trained within bounded error, but do not offer explicit control over the accuracy of their outputs. In this section, we examine the expressiveness of weak CoT models, incorporating inference-time strategies such as majority voting. The central question is: Can weak CoT be approximate samplers?
@@ 5.3.1 Formal Setting for Weak CoT
These questions are addressed in the subsequent sections, with formal proofs deferred to Appendix D and key ideas shown in Fig. 2.
@@ 5.3.3 Approximate Sampling via Self-Reducibility
Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation as defined in Definition C.4, and suppose further that $\sigma(x)=1$ for all $x\in\Sigma^{*}$. Let $p(y\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneqq\{y\mid(x,y)\in R\}$. Then, for CoT, the next-token prediction at each position $i$ corresponds to estimating the relative size of the subproblem
@@
Remark. The above result applies to any relation for which the CoT satisfies the weak token-wise learning assumption. However, the result becomes computationally meaningful and practically implementable when the relation is self-reducible, as shown in Proposition 5.12.

### v2

@@ Abstract
This paper is outdated.
For the latest version, please visit arXiv:2509.25239.
@@ 1 Introduction
  • Does separation exist beyond the logarithmic regime? → Affirmative; see Section˜4.
  • Are Looped Transformers universally more expressive than CoT? → Negative; see Section˜5.
@@ 2 Preliminaries and Notation
We consider models based on Transformer blocks that are applied recursively. Specifically, we analyze two architectures that differ in how intermediate computations are structured and reused across steps. We assume saturated attention, where attention weights are uniformly assigned to the tokens with the highest scores. For the feedforward computation, we consider either a standard fully connected layer or a mixture-of-experts (MoE) layer (Shazeer et al., 2017; Csordás et al., 2024). Formal definitions of the Transformer block are deferred to LABEL:app:tf.
@@ 2.3 Notation
For any $n\in\mathbb{N}^{+}$, we write $n]\coloneq\{1,2,\ldots,n\}$. For functions $f,g:\mathbb{N}\to\mathbb{N}$, we write $f(n)\in O(g(n))$ iff there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$. We denote by $\mathsf{poly}(n)$ the set of functions that grow at most polynomially, i.e., $\mathsf{poly}(n)\coloneq\left\{f:\mathbb{N}\to\mathbb{N}\;\middle|\;\exists k\in\mathbb{N},\;\exists c>0,\;\forall n\in\mathbb{N},\;f(n)\leq c\cdot n^{k}\right\}.$ Given two complexity classes ${\mathcal{C}}_{1}$ and ${\mathcal{C}}_{2}$, we write ${\mathcal{C}}_{1}\setminus{\mathcal{C}}_{2}$ to denote the set of decision problems (referred to as languages in computational complexity theory) that belong to ${\mathcal{C}}_{1}$ but not to ${\mathcal{C}}_{2}$.
@@ 3 General Problem Setting
This section begins by reviewing the problem settings considered in prior work, as summarized in Table˜1, and identifying the remaining open questions. We then formalize the two problem settings addressed in this study, specifying their underlying assumptions and objectives.
@@ 3.3 Our Problem Settings
We consider two fundamental classes of computational problems: the deterministic problem and the probabilistic problem, elaborated in Section˜4 and Section˜5, respectively. This section provides high-level definitions that specify the input–output structure in each setting; precise formulations are deferred to the corresponding sections. These problem classes differ in several fundamental aspects, such as output uniqueness, the presence of a target distribution, and the inclusion of accuracy parameters.
@@ 4 Looped Transformer Enables Efficient Parallel Solutions
This section analyzes the fundamental computational differences between CoT and Looped TF in the deterministic setting. We formalize deterministic computations using computation graphs, as introduced in Section˜4.1. In Section˜4.2, we show how CoT and Looped TF differ in processing the same computation graphs, highlighting their respective strengths. Finally, Section˜4.3 presents a separation in their computational complexity in the polylogarithmic regime.
@@ 4.1 Problem Setting
Let $\Sigma$ be a finite alphabet. A computation graph is a directed acyclic graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ that defines a function $G_{n}:\Sigma^{n}\to\Sigma^{m}$, where $\mathcal{F}_{n}=\{f_{n}:\Sigma^{\ell_{n}}\to\Sigma\mid\ell_{n}\in\mathbb{N}\}$ is a finite set of local functions associated with each node. The graph $G_{n}$ consists of (i) input nodes with in-degree 0, (ii) function nodes labeled by $f\in\mathcal{F}_{n}$ whose in-degree matches the arity of the corresponding function, and (iii) $m$ output nodes $O\subseteq V_{n}$ with out-degree 0. The overall function is obtained by evaluating the graph in topological order. The size of the graph is $|V_{n}|$, and its depth is the length of the longest path from an input node to an output node. We assume that the size of the computation graph $|V_{n}|$ is bounded by a polynomial in the input length $n$.
@@
To characterize the expressiveness of CoT and Looped TF, we analyze how each simulates computation graphs. We begin by formalizing the assumptions under which computations can be implemented by Transformer-based models (see LABEL:app:def for details).
@@
Remark. Proofs are given in LABEL:app:det, with illustrations in Fig. 1. Informally, Theorem˜4.4 shows that CoT simulates computations by decoding nodes in topological order, whereas Theorem˜4.7 shows that Looped TF simulates computations in a layer-wise and parallel fashion. This contrast reflects complementary strengths: CoT employs intermediate steps as scratchpad memory, supporting general computation without structural constraints. In contrast, Looped TF exploits structural parallelism to achieve greater efficiency, but requires additional assumptions.
@@ 4.3 Separation between CoT and Looped Transformers in Polylogarithmic Time
Let $\mathsf{CoT}[T(n),d(n),p(n)]$ (resp. $\mathsf{LOOP}[T(n),d(n),p(n)]$) denote the set of languages ${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a CoT (resp. Looped TF) $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(p(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token, after $O(T(n))$ decode steps (resp. loops), is ${\mathcal{L}}(x)$, i.e., $M_{n}(x)={\mathcal{L}}(x),\text{ for all }x\in\{0,1\}^{n}.$
@@
Since Boolean circuits allow polynomial-size, we assume a polynomial embedding size. Prior work has shown that CoT is equivalent to Boolean circuits composed of $\mathsf{TC}^{0}$ gates, with circuit “size” proportional to the number of CoT steps. This class $\mathsf{SIZE}^{\mathsf{TC}^{0}}$ is defined in LABEL:def:tcsize.
@@
There are several concrete problems in (nonuniform) $\mathsf{NC}^{2}$: solving linear equations, testing for perfect matching in bipartite graphs, and membership testing for fixed context-free grammars, denoted by LIN, MATCH, and FCFL, respectively (see LABEL:app:nc2 for details). Then we have:
@@ 5 Chain-of-Thought Enables Probabilistic Reasoning
In this section, we investigate the expressiveness of CoT with stochastic decoding for probabilistic tasks. Section˜5.1 introduces the task formulation and defines the model. Section˜5.2 analyzes the expressive power of CoT in an idealized setting and establishes a formal separation from Looped TF. Finally, Section˜5.3 shows that even weak CoT, when combined with self-consistency, exhibits sufficient expressive power.
@@ 5.1 Problem Setting
A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if any solution $y=y_{1}\dots y_{n}$ to input $x$ can be constructed incrementally, such that a prefix $y_{1:\sigma(x)}$ determines a subproblem $\psi(x,y_{1:\sigma(x)})$ whose solution completes $y$. See LABEL:def:self-red for a formal definition.
@@
Approximate sampling is closely related to approximate counting (Jerrum et al., 1986). For a relation $R$, define $f(x)\coloneq\lvert R(x)\rvert$, the number of solutions $y$ such that $(x,y)\in R$. A Fully Polynomial-Time Approximation Scheme (FPTAS) is a deterministic algorithm that, for any input $x$ and $\varepsilon>0$, outputs $\hat{f}(x)$ such that $(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ in time polynomial in $|x|$ and $1/\varepsilon$. A Fully Polynomial Randomized Approximation Scheme (FPRAS) is a randomized version that additionally takes $\delta>0$, and satisfies the same guarantee with probability at least $1-\delta$. See LABEL:app:approx for details.
@@ 5.3 Expressiveness of Weak CoT
While Lemma˜5.5 analyzes an idealized setting where a CoT, viewed as a PTM, is assumed to receive explicit error parameters, such assumptions diverge from the behavior of practical LLMs. In practice, models are trained within bounded error, but do not offer explicit control over the accuracy of their outputs. In this section, we examine the expressiveness of weak CoT models, incorporating inference-time strategies such as majority voting. The central question is: Can weak CoT be approximate samplers?
@@ 5.3.1 Formal Setting for Weak CoT
These questions are addressed in the subsequent sections, with formal proofs deferred to LABEL:app:prob and key ideas shown in Fig. 2.
@@ 5.3.3 Approximate Sampling via Self-Reducibility
Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation as defined in LABEL:def:self-red, and suppose further that $\sigma(x)=1$ for all $x\in\Sigma^{*}$. Let $p(y\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneqq\{y\mid(x,y)\in R\}$. Then, for CoT, the next-token prediction at each position $i$ corresponds to estimating the relative size of the subproblem
@@
Remark. The above result applies to any relation for which the CoT satisfies the weak token-wise learning assumption. However, the result becomes computationally meaningful and practically implementable when the relation is self-reducible, as shown in Proposition˜5.12.